Find the integral surface of the linear differential equation (x-y)y^2p+(y-x)x^2=(x^2+y^2)z, which passes through the curve xz=a^3,y=0

1
Expert's answer
2021-11-02T16:35:22-0400

"(x-y)y^2p+(y-x)x^2q=(x^2+y^2)z"


"\frac{dx}{(x-y)y^2}=\frac{dy}{(y-x)x^2}=\frac{dz}{(x^2+y^2)z}"


"x^2dx+y^2dy=0"

"x^3+y^3=c_1"


"\frac{dx-dy}{(x-y)(x^2+y^2)}=\frac{dz}{(x^2+y^2)z}"


"ln(x-y)=lnz+lnc_2"


"\frac{x-y}{z}=c_2"


"F(c_1,c_2)=F(x^3+y^3,\frac{x-y}{z})=0"


for  the curve "xz=a^3,y=0" :


"c_1=x^3"

"x\/z=c_2"

"\implies z=x/c_2=\sqrt[3]{c_1}/c_2"


"x/c_2=a^3/x\implies x^2=c_2a^3"


"z=\frac{z\sqrt[3]{x^3+y^3}}{x-y}"


"\sqrt[3]{x^3+y^3}=x-y"


"z=a^3/x=a^3/\sqrt[3]{c_1}=\frac{a^3}{\sqrt[3]{x^3+y^3}}"